d = c - s/n ( for s)
a = 1/2 h (B + b) (for h)
30.
Crafts. Rose’s garden is in the shape of a trapezoid. If the height of the trapezoid
is 16 m, one base is 20 m, and the area is 224 m2, find the length of the other base.

The sum of two consecutive integers is 145. Find the two
integers.
Crafts. Yuri has a board that is 98 in. long. He wishes to cut the board into two
pieces so that one piece will be 10 in. longer than the other. What should the length
of each piece be?

Answer by malakumar_kos@yahoo.com(315)

.



1) d = c-s/n or s/n = c-d therefore s = n(c-d) is the solution
2) a = 1/2.h(B+b) or 2a = h.(B+b) therefore h = 2a/(B+b) is the solution.
3)Height of the trapezoid h = 16 m one of the base = a = 20 m
area of the trapezoid A = 224 sqm to find the value of the other base = b
A = 1h(a+b)/2 2A = h(a+b) or 2A/h = (a+b) therefore (a+b) = 2.224/16 or a+b = 2.14 = 28 therefore a+b = 28 or b = 28-20 = 8 m
the value of the other base = b = 8m
4) Let the two consecutive numbers be x and (x+1) sum of the numbers = 145
therefore x+(x+1) = 145 i.e, 2x+1 = 145 , 2x = 145-1 = 144 or x=144/2
therefore x = 72 and x+1 = 72+1 = 73 the no's are 72 and73
5) length of the board = 98 inch let the length of the two pieces of boards
be x and (x+10)
therefore x+x+10 = 98 or 2x = 98-10 or 2x = 88 x = 88/2 = 44
the length of one board = x = 44 inch
the length of the other board = x+10 =44+10 =54 inch